Use an appropriate method to solve the given system e *-=( )x+ te

please show work and steps thank you
Answer #1

Given that, x 8 xt (e e-t tet Let A= 18 I -1 The characteristic equation is, A-NI co 1-X o 8 -1- : (t-> )(-1-))-8=0 72-9=0 --det+ b[tet tet = [ 24+ (!)] ett [bAt(o)tet Comparing the coefficients we get a At b = la Aten Let as coun) bellen Now fa 8 +(E)(3)(..)+10 ( ( but $b2 ) + (0) bi-bz p = b 1-351 b = bi-b2+1 8.62= bi+1=0 bit b=0 (bs) o xp = et elis) + tet (o) ) + tet 8